# A Practical Probabilistic Benchmark for AI Weather Models

### v1

@@ 3.2 Comparing data-driven and NWP models with the same ensemble method
Figure 3 compares the ensemble metrics for a set of forecast models including GraphCast, Pangu Weather, and IFS HRES (see Sec. A.3 for details). From the vantage of deterministic skill scores (left column) both of these AI weather models appear to outperform the IFS, and GraphCast especially and increasingly so as a function of lead time. However, the ensemble skill scores – CRPS and RMSE of the ensemble mean – tell a different story. All models are virtually tied, subtle differences between the models’ ensemble RMSE shrink rather than grow with lead time, and (bottom right) Pangu Weather actually has the lowest Z500 CRPS for many lead times. Overall, these results demonstrate a non-monotonic relationship between deterministic skill and the performance of an ensemble. Also we have shown that some data-driven models, despite being optimized on deterministic loss functions, can outperform physics-based models such as IFS when the same ensemble technique is used.
@@ 3.3 The effect of long-lead time training
LEF confirms this hypothesis. We begin by training SFNO models with varying amounts of long-lead time fine-tuning of the autoregressive training. The base model is trained for 80 epochs with a context length of a single 6-hour timestep. We also include results with 2, 3, 5, 9, and 11 steps of autoregressive timestepping in the MSE loss (12h-3 days) to illustrate the effect. The pre-trained base model is fine-tuned with a smaller learning rate for 20 epochs with a context length of two 6-hour timesteps. The output from a single forward pass serves as the input for the model to generate a 12-hour forecast. The total accumulated loss from the 6 and 12-hour forecasts is then used as the training objective. This fine-tuning procedure is sequentially extended, with the context length progressively increasing from 18 hours (3 timesteps) to 3 days (11 timesteps). For each increment in context length, the model is trained for 6 epochs. NVIDIA has previously released the 11-step model with the name “FourCastNet v2-small”  [^2] This model is available on NVIDIA’s NGC Catalog and also on ECMWF’s ai-models platform.
@@
Figure 4: Using LEF to explore the effect of fine-tuning. Deterministic RMSE (a), ensemble mean RMSE (b) and CRPS (c) as a function of lead time. Ensemble standard deviation as a function of ensemble mean RMSE (d). The color lines show a range of fine tuning steps and the dashed line shows the ifs. The field being scored is temperature at 850 hPa (T850).
@@ 3.4 Sensitivity to effective resolution
We now show that it is possible to make a data-driven forecast more dispersive by tuning the architecture. For computational reasons, many ML models (see Sec. 1) project the ${}^{\circ}$ inputs into a coarse latent space early in the network architecture. We suspect the resolution of this latent space limits the resolutions the architecture can faithfully generate in autoregressive forecast.
@@
Figure 5: Sensitivity of the ensemble spread to changes in hyperparameters. (a-d) spread/skill results for various fields and differing amounts of auto-regressive fine tuning. (e-h) shows the same, but with a varying scale factor.
@@ 4 Discussion and Conclusions
On the other hand, much of the data-driven forecast literature has focused primarily on minimizing the error in deterministic forecasts. The easiest way to do this is to reduce the variance of a forecast. It is easy to recognize ensemble averaging or space-time filtering as variance-reduction techniques. However, there are a large number of other ML design choices whose effects on ensemble dispersion are less obvious. For instance, we have identified that effective resolution can be a helpful lever on dispersion separately from determinstic skill.
In this work, we devise a parsimonious framework that allows one to compare physics-based and data-driven models with the same ensemble technique. Within this framework, we highlight that the choice of loss function can produce artificial dissipation akin to ensemble averaging. Many recent papers have used some form of training involving multi-day lead times including, for example, auto-regressive MSE losses(Lam et al., 2023; Pathak et al., 2022), replay buffers (Chen et al., 2023a, b), or simply predicting very long lead times directly Nguyen et al. (2023). Our results show that, when taken too far, such loss functions lead to a poorly calibrated ensemble.
While lagged ensemble scoring is a practical framework for comparing deterministic forecasts, several improvements are possible. First of all, lagged ensemble members are not interchangeable so we know that a weighted ensemble will outperform our unweighted ensembles here. The analysis of ensemble calibration is made more ambiguous by this subtlety; we haved use the LEF of IFS as a baseline to circumvent this issue. Future work can also focus on applying this technique to more models and working to integrate it into existing community benchmarks which provide a uniform protocol for e.g. output resolution (Rasp et al., 2023b).
Finally, this work tests a deterministic forecast in isolation from other components. This is most useful if we assume that the probabilistic forecast systems of the future will be built as they are today: from ensembles of deterministic time steppers, but recent work on intrinsically stochastic models suggests an alternative path (Price et al., 2023; Kochkov et al., 2023). While we could score such models with lagged ensembles, this benchmark may give an unfair advantage to deterministic models that have not yet solved the more complex end-to-end probabilistic prediction task.
@@ Acknowledgments
Appendix A Methods
@@ A.2 Numerical models
The European Centre Medium Range Forecast’s (ECMWF) hosts the world’s leading global weather forecasting systems. Its deterministic model, the High RESolution forecast (HRES) is a configuration of the Integrated Forecast System (IFS) which produces global 10-day forecasts at about $0.1^{\circ}$ resolution ECMWF [2019b]. It is used operationally as the high resolution deterministic member of an ensemble forecasting system (ENS) where additional 50 members at $0.25^{\circ}$ are used to generate the operational forecast.
@@ A.3 Data-driven models
Many models are trained with autoregressive loss accumulation of different kinds (see Table 1). Brenowitz and Bretherton [2018] introduced this idea to stabilize machine learning parameterizations, and then Weyn et al. [2021] applied it to global weather prediction. Since then, severals works have also used autoregressive loss, including SFNO [Bonev et al., 2023] and GraphCast [Lam et al., 2023]. The latter reported a comparison of the effect of number of steps on the RMSE (see figure S34 there) and also note that auto-regressive training encourages GraphCast to blur the forecast. PanguWeather [Bi et al., 2023] took a different approach by training several models at different time-steps (namely 1,3,6,24 hours) and interleaving them in the forecast. FuXi Chen et al. [2023b] combines both methods in the so-called cascade approach where separate models are trained and fine-tuned for different parts of the medium range (0-14 days lead time). FengWu [Chen et al., 2023a] used a “replay buffer” in which the model is trained to predict the weather at time $t$ from both target at time $t-1$ and from stored model forecasts at time $t-1$; effectively learning to veer back to the target data when provided with initial condition with a lead time error accumulated over several steps. This in practice could encourage the model to learn an approximation of the ensemble mean as well.
Furthermore, most models are evaluated with deterministic metrics (RMSE and ACC), exceptions are reported in Bi et al. [2023], Chen et al. [2023a].
Table 1: A list of the most relevant global ML weather models. Here NLL used in Chen et al. [2023a] is a negative log-likelihood loss.
@@
Architecture
CNN
SFNO
GNN
swin
Transformer
Enc-Dec
Transformer
@@
$2^{\circ}$
$0.25^{\circ}$
$0.25^{\circ}$
$0.25^{\circ}$
$0.25^{\circ}$
$1^{\circ}$
$0.25^{\circ}$
@@
4 $\times$ 6h steps
12 $\times$ 6h steps
12 $\times$ 6h steps
interleaving 1,3,6,24h
12 $\times$ 6h steps + interleaving
10 $\times$ 6h steps
Replay Buffer
@@ 10 $\times$ 6h steps
RMSE, ACC
RMSE, ACC
RMSE, ACC
RMSE, ACC, CRPS, STD/RMSE
RMSE, ACC, CRPS, STD, SSR
RMSE
ACC, RMSE
@@
Bonev et al. [2023] introduced a spherical Neural Operator, which respects the underlying geometry of the sphere and its inherent symmetries. The resulting architecture has remarkable stability properties with reported autoregressive rollouts of up to a full year Bonev et al. [2023].
@@
Pangu Weather Bi et al. [2023] introduced a hierarchy of vision transformers trained to take different time steps. The publicly available models include
@@
The interleaved time stepper is effectively a 24-hour model, with 6-hour steps used to interleave in between. For example, to make a 72 hour forecast, requires 3 24 hour steps and no 6 hour steps. A 36 hour forecast requires a single 24 time step and two 6 hour steps.
@@ A.4 Ensemble Verification Metrics
Probabilistic predictions aim to maximize sharpness subject to calibration Raftery et al. [2005]. Qualitatively, calibration means that the likelihood of observing the true value is the same as observing a member drawn from the ensemble. A necessary condition for calibration is that the spread-error relationship be 1-to-1 when averaged over a sufficient sample [Fortin et al., 2014]. With this condition in hand, reducing RMSE of the ensemble mean is the same as reducing the ensemble spread.
The calibrate-sharpness perspective suggests that ensemble-mean RMSE is equivalent to the spread of an ensemble calibrated by variance inflation. In the limit of many ensemble members artificially increasing the spread won’t change the ensemble mean much, since the variance of the ensemble mean is given by $\sigma^{2}/n=o(1/n)$. Therefore the spread (i.e. sharpness) of the calibrated ensemble is the same as the ensemble-mean RMSE. Thus, the ensemble with the lowest ensemble-mean RMSE regardless of spread is the better ensemble when viewed from the calibrate-sharpness perspective.
On the other hand, not all ensembles are calibrated. For these, we can compare the performance using proper scoring rules like continuous-ranked-probability score (CRPS)[Gneiting and Raftery, 2007]. CRPS is given by
@@
where $H(x)$ is the heaviside function and $F(x)=n^{-1}\sum_{k}H(x-x_{k})$ is the empirical cumulative distribution function, and $y$ is the observation.
We present area-weighted global averages of this metric, where the area weighting is given by $\cos(lat)$. Unlike Rasp et al. [2023b], for RMSE and spread, initial time averages are taken across the lagged ensemble at fixed lead time before the square root, so that the square of RMSE is a time and space-averaged MSE.

### v2

@@ 3.2 Comparing data-driven and NWP models with the same ensemble method
Figure 2 compares the ensemble metrics for a set of forecast models including GraphCast, Pangu Weather, and IFS HRES (see Sec. A.3 for details). From the vantage of deterministic skill scores (left column) both of these AI weather models appear to outperform the IFS, and GraphCast especially and increasingly so as a function of lead time. However, the ensemble skill scores – CRPS and RMSE of the ensemble mean – tell a different story. All models are virtually tied, subtle differences between the models’ ensemble RMSE shrink rather than grow with lead time, and (bottom right) Pangu Weather actually has the lowest Z500 CRPS for many lead times. The relative performance of Graphcast and Pangu with respect to IFS HRES is similar for different ensemble sizes and channels (see Figure A2). Overall, these results demonstrate a non-monotonic relationship between deterministic skill and the performance of an ensemble. Also we have shown that some data-driven models, despite being optimized on deterministic loss functions, can outperform physics-based models such as IFS when the same ensemble technique is used.
@@ 3.3 The effect of long-lead time training
LEF confirms this hypothesis. We begin by training SFNO models with varying amounts of long-lead time fine-tuning of the autoregressive training. The base model is trained for 80 epochs with a context length of a single 6-hour timestep. We also include results with 2, 3, 5, 9, and 11 steps of autoregressive timestepping in the MSE loss (12h-3 days) to illustrate the effect. The pre-trained base model is fine-tuned with a smaller learning rate for 20 epochs with a context length of two 6-hour timesteps. The output from a single forward pass serves as the input for the model to generate a 12-hour forecast. The total accumulated loss from the 6 and 12-hour forecasts is then used as the training objective. This fine-tuning procedure is sequentially extended, with the context length progressively increasing from 18 hours (3 timesteps) to 3 days (11 timesteps). For each increment in context length, the model is trained for 6 epochs. NVIDIA has previously released the 11-step model with the name “FourCastNet v2-small”  [^1] This model is available on NVIDIA’s NGC Catalog and also on ECMWF’s ai-models platform..
@@
Figure 3: Using LEF to explore the effect of fine-tuning. Deterministic RMSE (a), ensemble mean RMSE (b) and CRPS (c) as a function of lead time. Spread as a function of ensemble mean RMSE (d). The spread is multiplied by a factor involving the ensemble size $R=2M+1=9$ (see Appendix A.4). The color lines show a range of fine tuning steps and the dashed line shows the ifs. The field being scored is temperature at 850 hPa (T850). Unlike Figure 2 the typical biased estimator of CRPS is used (5).
@@ 3.4 Sensitivity to effective resolution
We now show that it is possible to make a data-driven forecast more dispersive by tuning the architecture. For computational reasons, many ML models (see Sec. 1) project the 0.25 inputs into a coarse latent space early in the network architecture. We suspect the resolution of this latent space limits the resolutions the architecture can faithfully generate in autoregressive forecast.
@@
Figure 4: Sensitivity of the ensemble spread to changes in hyperparameters. (a-d) show results for various fields and differing amounts of auto-regressive fine tuning. (e-h) shows the same, but with a varying scale factor. The bias-corrected spread error ratio is defined by $\text{SER}=\sqrt{\frac{R+1}{R}}\frac{\text{Spread}}{\text{eRMSE}}$.
@@ 4 Discussion and Conclusions
On the other hand, while there has been substantial progress towards probabilistic forecast (Price et al., 2023; Kochkov et al., 2023), much of the data-driven forecast literature has focused primarily on minimizing the error in deterministic forecasts The easiest way to do this is to reduce the variance of a forecast. It is easy to recognize ensemble averaging or space-time filtering as variance-reduction techniques. However, there are many other ML design choices whose effects on ensemble dispersion are less obvious. For instance, we have identified that effective resolution can be a helpful lever on dispersion separately from determinstic skill.
In this work, we devise a simple framework that allows one to compare physics-based and data-driven models with the same ensemble technique. Within this framework, we highlight that the choice of loss function can produce artificial dissipation akin to ensemble averaging. Many recent papers have used some form of training involving multi-day lead times including, for example, auto-regressive MSE losses(Lam et al., 2023; Pathak et al., 2022), replay buffers (Chen et al., 2023a, b), or simply predicting very long lead times directly (Nguyen et al., 2023). Our results show that, when taken too far, such loss functions lead to a poorly calibrated ensemble.
While lagged ensemble scoring is a practical framework for comparing deterministic forecasts, several improvements are possible. First of all, lagged ensemble members are not interchangeable so we know that a weighted ensemble will outperform our unweighted ensembles here. The analysis of ensemble calibration is made more ambiguous by this subtlety; we have used the LEF of IFS as a baseline to circumvent this issue. Finally ensemble size of a lagged ensemble is inherently limited especially if we want the ensemble to be meaningful. Future work can also focus on applying this technique to more models and working to integrate it into existing community benchmarks which provide a uniform protocol for e.g. output resolution (Rasp et al., 2023).
Finally, this work presents a convenient way to test a deterministic forecast in isolation from other components. This is most useful if we assume that the probabilistic forecast systems of the future will be built as they are today: from ensembles of deterministic time steppers. Indeed AI weather models in this category have been shown to be capable of competitive probabilistic prediction skill (Mahesh et al., 2024). However, recent work on intrinsically stochastic models suggests an alternative path (Price et al., 2023; Kochkov et al., 2023). While we could score such models with lagged ensembles, this benchmark may give an unfair advantage to deterministic models that have not yet solved the more complex end-to-end probabilistic prediction task. So final evaluations of end-to-end ensemble forecast systems should still be done with traditional ensemble scoring methods (Rasp et al., 2023). Future work could explore more convenient scoring protocols of end-to-end systems—for example with very small ensembles and unbiased scoring rules.
@@ Acknowledgments
Open Research
Software
A parallel implementation of the lagged ensemble scoring and model inference capability is available at http://github.com/NVIDIA/earth2mip.
Data
The ERA5 dataset used for training and evaluation is widely available. We obtained it from a mirror of the Research Data Archive at NCAR European Centre for Medium-Range Weather Forecasts (2017) stored at NERSC. The IFS ensemble and HRES hindcasts that we used as baselines were downloaded from the TIGGE archive at ECWMF.
Model Weights
The graphcast model weights were downloaded from a Google cloud storage bucket gs://dm_graphcast. The PanguWeather weights were downloaded from the ECWMF AI Models repository. The 11-step trained SFNO checkpoint (cf. Fig 3) is available at https://catalog.ngc.nvidia.com/orgs/nvidia/teams/modulus/models/modulus_fcnv2_sm. The earth2mip software automatically downloads these files and can be examined to see the precise URLs used.
Appendix A Supplemental Information
@@ A.2 Numerical models
The European Centre Medium Range Forecast’s (ECMWF) hosts the world’s leading global weather forecasting systems. Its deterministic model, the High RESolution forecast (HRES) is a configuration of the Integrated Forecast System (IFS) which produces global 10-day forecasts at about $0.1^{\circ}$ resolution [ECMWF, 2019b]. It is used operationally as the high resolution deterministic member of an ensemble forecasting system (ENS) where additional 50 members at $0.25^{\circ}$ are used to generate the operational forecast.
@@ A.3 Data-driven models
Many models are trained with autoregressive loss accumulation of different kinds (see Table 1). Brenowitz and Bretherton [2018] introduced this idea to stabilize machine learning parameterizations, and then Weyn et al. [2021] applied it to global weather prediction. Since then, severals works have also used autoregressive loss, including SFNO [Bonev et al., 2023] and GraphCast [Lam et al., 2023]. The latter reported a comparison of the effect of number of steps on the RMSE (see figure S34 there) and also note that auto-regressive training encourages GraphCast to blur the forecast. PanguWeather [Bi et al., 2023] took a different approach by training several models at different time-steps (namely 1,3,6,24 hours) and interleaving them in the forecast. FuXi [Chen et al., 2023b] combines both methods in the so-called cascade approach where separate models are trained and fine-tuned for different parts of the medium range (0-14 days lead time). FengWu [Chen et al., 2023a] used a “replay buffer” in which the model is trained to predict the weather at time $t$ from both target at time $t-1$ and from stored model forecasts at time $t-1$; effectively learning to veer back to the target data when provided with initial condition with a lead time error accumulated over several steps. This in practice could encourage the model to learn an approximation of the ensemble mean as well.
Furthermore, most models are evaluated with deterministic metrics (RMSE and ACC), exceptions are reported [Bi et al., 2023, Chen et al., 2023a].
Table 1: A list of the most relevant global ML weather models. Here NLL used by Chen et al. [2023a] is a negative log-likelihood loss.
@@
Architecture CNN SFNO GNN swin Transformer Enc-Dec Transformer
@@
Bonev et al. [2023] introduced a spherical Neural Operator, which respects the underlying geometry of the sphere and its inherent symmetries. The resulting architecture has remarkable stability properties with reported autoregressive rollouts of up to a full year [Bonev et al., 2023].
@@
Pangu Weather [Bi et al., 2023] introduced a hierarchy of vision transformers trained to take different time steps. The publicly available models include
@@
The interleaved time stepper is effectively a 24-hour model, with 6-hour steps used to interleave in between. For example, to make a 72 hour forecast, requires 3 24 hour steps and no 6 hour steps. A 36 hour forecast requires a single 24 time step and two 6 hour steps. We use the same interleaving strategy for each member of the lagged ensemble, so that a -1 lagged member at time $t$ is the same as the 0-lagged member at time $t-1$. Each lagged member is simply a deterministic forecast at the lagged initial time.
@@ A.4 Ensemble Verification Metrics
The deterministic RMSE is given by
$\text{dRMSE}=\sqrt{\langle|x_{0}-y|^{2}\rangle}.$ (2)
When evaluating with a small ensemble size some care is required in defining and interpreting probabilistic metrics. Let the ensemble mean $\overline{x}=R^{-1}\sum_{m=1}^{R}x_{m}$. We use the common definition of ensemble mean MSE:
$\text{eRMSE}=\sqrt{\left\langle\left|\overline{x}-y\right|^{2}\right\rangle}$ (3)
This formula is biased for finite ensembles. Zamo and Naveau [2018] give an unbiased estimator of CRPS as
$\text{CRPS}=\left\langle\frac{1}{R}\sum_{m=1}^{R}\left|x_{m}-y\right|-\frac{1}{2R(R-1)}\sum_{m,m^{\prime}=1}^{R}\left|x_{m}-x_{m}^{\prime}\right|\right\rangle.$ (4)
The more commonly used, yet biaesd formula—derived from estimating using the substituting the empirical CDF into the continuous formula—is given by
$\text{CRPS}_{biased}=\left\langle\frac{1}{R}\sum_{m=1}^{R}\left|x_{m}-y\right|-\frac{1}{2R^{2}}\sum_{m,m^{\prime}=1}^{R}\left|x_{m}-x_{m}^{\prime}\right|\right\rangle.$ (5)
Note the different denominator in the second term leads to $\text{CRPS}_{biased}$ favoring models with lower spread.
The ensemble spread is defined as the unbiased sample standard deviation of the ensemble, which is given by
$\text{Spread}=\sqrt{\left\langle\frac{1}{R-1}\sum_{m=1}^{R}\left|x_{m}-\overline{x}\right|^{2}\right\rangle}.$ (6)
Fortin et al. [2014] show that that a necessary condition for a well calibrated is that a de-biased spread-error ratio (SER) is 1. That is
$\text{SER}:=\sqrt{\frac{R+1}{R}}\frac{\text{Spread}}{\text{eRMSE}}=1.$ (7)
@@ A.5 Spread-error correspondence for a weighted ensemble
A.6 Sensitivity to Ensemble Size
M=1 M=2 M=3 M=4
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure A1. Same as 1b but for differing numbers of lags $M=1,2,3,4$ and for two channels, u10m and z500. The correlation between the lagged and CRPS is similar in all cases. The lagged CRPS has similar quantitative values to the ENS CRPS for $M=1$
Figure A2. Sensitivity of relative performance to the lagged ensemble size $M=1,4$. Shows the fractional improvement in CRPS relative to HRES for PanguWeather and Graphcast for several channels.